OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) x + 5y = 3 2x + 9y = 4 Okay, i have tried this about 10 times now. i can't get it i keep coming out to 4,16, and 34.

OpenStudy (anonymous):

& i have to do this!???! for the next proble? Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) x + 4y = –16 3x + 2y = 12

OpenStudy (mertsj):

How can you get three answers for two variables?

OpenStudy (anonymous):

so you see my problem?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

multiply the first equation by -2, add to second one

OpenStudy (anonymous):

oh wait, no the first times i got 4 & 16 than i got 16 and 34 and vise versa? but no other numbers?

OpenStudy (anonymous):

\[-2x-10y=-6\] \[2x+9y=4\] \[-y=-2\] \[y=2\]

OpenStudy (anonymous):

how?!

OpenStudy (anonymous):

so solution to \[x + 5y = 3, 2x + 9y = 4 \] is \[x=-7,y=2\]

OpenStudy (anonymous):

okkk okk hold on let me write that down~~~~~

OpenStudy (anonymous):

okay! got it

OpenStudy (mertsj):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) To solve by substitution, solve the first equation for x and replace x in the second equation with that expression. The resulting equation will have only one variable and it will be x. Solve that equation for x. Replace x in either equation with that value and solve for y.

OpenStudy (mertsj):

Satellite did Part !!

OpenStudy (anonymous):

okay is it weird so say ilove you both?

OpenStudy (mertsj):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) First, solve each equation for y so that both equations are in the slope-intercept form, y = mx +b where m is the slope and b is the y intercept. Mark the number b on the y axis and then follow the directions given by the slope which is rise/run. Then draw the line. Repeat with the second equation.

OpenStudy (mertsj):

Part 3: Explain what the solution represents on the graph. (1 point) The solution is the point where the lines cross on the graph.

OpenStudy (mertsj):

No. It is not weird. We love you too.

OpenStudy (anonymous):

lol awesome!!! :)))